# GUAVA: Generalizable Upper Body 3D Gaussian Avatar

### v1

@@ Abstract
Figure 1: For each single image, GUAVA can reconstruct a 3D upper-body Gaussian avatar via feed-forward inference within sub-second time, enabling real-time expressive animation and novel view synthesis at 512$\times$512 resolution.
@@ 2 Related work
Traditional human or head reconstruction primarily focuses on mesh reconstruction. Researchers have constructed human template models like SMPL [55], the facial model FLAME [46], and the hand model MANO [78] based on thousands of 3D scans. These models are mesh-based and represent shape and expression variations in a linear space, with rotational movement modeled using joints and linear blend skinning. SMPLX [65] extends SMPL by integrating FLAME and MANO, enhancing facial and hand expressiveness. Since SMPLX is trained on full-body scans and may overlook facial details, its expressive capability for face is still limited compared to FLAME. By combining deep learning, some methods [28, 103, 80, 10, 41, 21, 25, 15, 75] predict the parameters of human, hand, and face models to achieve fast image-to-mesh reconstruction.
With the development of Neural Radiance Fields (NeRF) [59], studies have combined NeRF with template models for more realistic 3D reconstructions of heads [109, 1, 73] or humans [45, 95, 90]. NeRFcae [26] extends NeRF to a dynamic form by introducing expression and pose parameters as conditioning inputs for driveable avatar reconstruction. Neural Body [69] proposes a new representation for dynamic humans by encoding posed human meshes into latent code volumes. Recently, 3DGS [40] has made breakthroughs in real-time rendering, leading to its rapid application in various fields [72, 105, 27, 36, 88], leveraging its efficient rendering capabilities, including avatar reconstruction [34, 44, 57, 79, 52, 104]. GART [44] uses 3D Gaussians to represent deformable subjects, employing learnable blend skinning to model non-rigid deformations and generalizing to more complex deformations with novel latent bones. GaussianAvatar [33] learns pose-dependent Gaussians using a 2D pose encoder to represent human avatar. ExAvatar [61] combines mesh and Gaussian representations, treating each Gaussian as a vertex in the mesh with pre-defined connectivity, enhancing facial expression expressiveness.
@@
Figure 2: Given the source and target images, we first obtain the shape, expression, and pose parameters of the EHM template model through tracking. The source image is then passed through an image encoder to extract an appearance feature map. Using these features and the tracked EHM, one branch predicts the template Gaussians, and the other predicts the UV Gaussians. These are combined to form the Ubody Gaussians in canonical space, which are then deformed into pose space using the tracked parameters from the target image. Finally, a coarse feature map is rendered and refined by a neural refiner to produce the final image.
@@ 3.2 Reconstruction Modeling
UV Gaussians. Using only the EHM’s vertices to represent the avatar can not fully capture high-frequency details due to the limited number of vertices, and struggles to represent regions beyond the template model. To address these issues, we propose introducing additional Gaussians to capture fine details. Specifically, we use the UV texture map to construct UV Gaussians, where we predict a Gaussian for each valid pixel in the texture map. Unlike template Gaussians, which are directly represented in world coordinates, as in [70] we rig each UV Gaussian to the corresponding mesh triangle. Each UV Gaussian is defined as $G_{uv}=\{\Delta{\mu},r,s,\alpha,c,k,t\}$, where $\Delta{\mu}$ represents the local position of the Gaussian in the triangle, and $t$ refers to interpolated position on the triangle using the pixel’s barycentric coordinates. During rendering, these properties are transformed into world coordinates:
@@ 4.1 Experimental Setup
Implementation details. Our model is built with PyTorch [64] and trained using the Adam optimizer [43]. We train the model on NVIDIA RTX A6000 GPUs for 200,000 iterations, consuming about 78 GPU hours, with a total batch size of 12. For data processing, we use PIXIE [24], TEASER [54], and HaMeR [66] to perform rough estimations of SMPLX, FLAME, and MANO parameters, respectively. Body masks are extracted using [11], and the background is set to black. Moreover, we extract human and facial keypoints using [101, 56, 4, 16].
@@ 4.2 Evaluation
Quantitative results. Self-reenactment. For 2D-based methods, we evaluate all video frames, while for 3D-based methods, we assess the latter half. As shown in Tab. 1 and Tab. 2, our method outperforms all others across all metrics, demonstrating its ability to reconstruct high-quality avatars with accurate motion and photorealistic rendering. Cross-reenactment. We use 10 source images from the test set and ViCo-X [112], driven by 8 videos (5079 frames). Tab. 3 shows that our method significantly outperforms others in IPS, proving its superior ability to preserve identity consistency with the source ID under different poses. Efficiency. We evaluate different methods on the NVIDIA RTX 3090 GPU, with results shown in Tab. 1 and Tab. 2. Compared to 2D-based methods, our approach achieves around 50 FPS in animation and rendering, while others take several seconds per frame. 3D-based methods also support real-time rendering but require minutes to hours for reconstruction, whereas our method completes it in just 0.1 seconds.
@@ C.1 Novel views synthesis
C.2 EHM tracking results
Although we have demonstrated the improvement in model performance with the EHM model from both qualitative and quantitative perspectives in the main paper Sec. 4.3, to provide a more intuitive comparison, we further present the visual tracking results in Fig. 9. It is important to note that "w/o EHM" refers to tracking with SMPLX, which still uses our designed tracking framework, but without the FLAME integration step. From the results, it is clear that SMPLX can only roughly capture mouth movements, while EHM captures detailed facial expressions. Furthermore, our designed tracking framework not only captures accurate facial expressions but also tracks fine gestures, including finger movements, with high precision.
Appendix D Ethical considerations

### v2

@@ Abstract
Figure 1: From a single image with a tracked pose, GUAVA can reconstruct a 3D upper-body Gaussian avatar via feed-forward inference within sub-second time, enabling real-time expressive animation and novel view synthesis at 512$\times$512 resolution.
@@ 2 Related work
Traditional human or head reconstruction primarily focuses on mesh reconstruction. Researchers have constructed human template models like SMPL [57], the facial model FLAME [48], and the hand model MANO [79] based on thousands of 3D scans. These models are mesh-based and represent shape and expression variations in a linear space, with rotational movement modeled using joints and linear blend skinning. SMPLX [66] extends SMPL by integrating FLAME and MANO, enhancing facial and hand expressiveness. Since SMPLX is trained on full-body scans and may overlook facial details, its expressive capability for face is still limited compared to FLAME. By combining deep learning, some methods [30, 103, 81, 12, 43, 23, 27, 17, 109, 76] predict the parameters of human, hand, and face models to achieve fast image-to-mesh reconstruction.
With the development of Neural Radiance Fields (NeRF) [60], studies have combined NeRF with template models for more realistic 3D reconstructions of heads [111, 2, 74, 1] or humans [47, 95, 90]. NeRFcae [28] extends NeRF to a dynamic form by introducing expression and pose parameters as conditioning inputs for driveable avatar reconstruction. Neural Body [70] proposes a new representation for dynamic humans by encoding posed human meshes into latent code volumes. Recently, 3DGS [42] has made breakthroughs in real-time rendering, leading to its rapid application in various fields [73, 105, 29, 38, 11], leveraging its efficient rendering capabilities, including avatar reconstruction [36, 71, 80, 54, 104, 106]. GART [46] uses 3D Gaussians to represent deformable subjects, employing learnable blend skinning to model non-rigid deformations and generalizing to more complex deformations with novel latent bones. GaussianAvatar [35] learns pose-dependent Gaussians using a 2D pose encoder to represent human avatar. ExAvatar [62] combines mesh and Gaussian representations, treating each Gaussian as a vertex in the mesh with pre-defined connectivity, enhancing facial expression expressiveness.
@@
Figure 2: Given the source and target images, we first obtain the shape, expression, and pose parameters of the EHM template model through preprocessing tracking. The source image is then passed through an image encoder to extract an appearance feature map. Using these features and the tracked EHM, one branch predicts the template Gaussians, and the other predicts the UV Gaussians. These are concatenated to form the Ubody Gaussians in canonical space, which are then deformed into pose space using the tracked parameters from the target image. Finally, a coarse feature map is rendered and refined by a neural refiner to produce the final image.
@@ 3.2 Reconstruction Modeling
UV Gaussians. Using only the EHM’s vertices to represent the avatar can not fully capture high-frequency details due to the limited number of vertices, and struggles to represent regions beyond the template model. To address these issues, we propose introducing additional Gaussians to capture fine details. Specifically, we use the UV texture map to construct UV Gaussians, where we predict a Gaussian for each valid pixel in the texture map. Unlike template Gaussians, which are directly represented in world coordinates, as in [71], we rig each UV Gaussian to the corresponding mesh triangle. Each UV Gaussian is defined as $G_{uv}=\{\Delta{\mu},r,s,\alpha,c,k,t\}$, where $\Delta{\mu}$ represents the local position of the Gaussian in the triangle, $k$ is the triangle index, and $t$ refers to interpolated position on the triangle using the pixel’s barycentric coordinates. During rendering, these properties are transformed into world coordinates:
@@ 4.1 Experimental Setup
Implementation details. Our model is built with PyTorch [65] and trained using the Adam optimizer [45]. We train the model on NVIDIA RTX A6000 GPUs for 200,000 iterations, consuming about 156 GPU hours, with a total batch size of 12. For data processing, we use PIXIE [26], TEASER [56], and HaMeR [67] to perform rough estimations of SMPLX, FLAME, and MANO parameters, respectively. Body masks are extracted using [13], and the background is set to black. Moreover, we extract human and facial keypoints using [101, 58, 5, 18].
@@ 4.2 Evaluation
Quantitative results. Self-reenactment. For 2D-based methods, we evaluate all video frames, while for 3D-based methods, we assess the latter half. As shown in Tab. 1 and Tab. 2, our method outperforms all others across all metrics, demonstrating high-quality avatar reconstruction with accurate motion and photorealistic rendering. Cross-reenactment. We use 10 source images from the test set and ViCo-X [114], driven by 8 videos (5079 frames). Tab. 3 shows that our method significantly outperforms others in IPS, proving its superior ability to preserve identity consistency with the source ID under different poses. Efficiency. We evaluate all methods on an NVIDIA RTX 3090 GPU, with results shown in Tab. 1 and Tab. 2. Our method achieves around 50 FPS in animation and rendering, while other 2D-based methods take several seconds per frame. 3D-based methods also support real-time rendering but require minutes to hours for reconstruction, whereas our method completes it from a tracked image in just 0.1s.
@@ B.4 Inverse texture mapping
Figure 9: Visual results on extrapolated novel view synthesis. Renderings of back-facing regions are slightly lower quality due to the lack of backside data in our training set.
Figure 10: Full-Body reconstruction results. Our method successfully performs full-body reconstruction. Acquiring additional full-body data is expected to improve the results in such cases.
Figure 11: Visualization of failure cases. Our method exhibits limitations when handling fluffy hair, loose clothing, and flowing folds.
@@ C.1 Novel views synthesis
We also demonstrate novel view synthesis from extreme angles, as shown in Fig. 9. Since our training dataset includes only front views, back rendering is suboptimal. Training on 360° datasets, as IDOL [116] could help.
C.2 Full-body
Our method is also applicable to full-body settings as Fig. 10. However, due to few full-body data in our dataset, adding more data could improve generalization.
C.3 Failure cases
As shown in Fig. 11, our method has limitations with fluffy hair, loose clothing, and flowing folds — all calling for future improvements.
C.4 EHM tracking results
Although we have demonstrated the improvement in model performance with the EHM model from both qualitative and quantitative perspectives in the main paper Sec. 4.3, to provide a more intuitive comparison, we further present the visual tracking results in Fig. 12. It is important to note that ”w/o EHM” refers to tracking with SMPLX, which still uses our designed tracking framework, but without the FLAME integration step. From the results, it is clear that SMPLX can only roughly capture mouth movements, while EHM captures detailed facial expressions. Furthermore, our designed tracking framework not only captures accurate facial expressions but also tracks fine gestures, including finger movements, with high precision.
Appendix D More discussion
D.1 EHM vs SMPLX
As discussed in the main paper Sec. 2.1, although SMPLX [66] integrates SMPL [57] and FLAME [48], its expression space is newly trained on full-body scans, which may overlook fine facial details. This results in SMPLX having less expressive facial expressiveness compared to FLAME, a limitation also noted in ExAvatar (Sec. 3.1) [62]. EHM’s main contribution is to improve this facial expressiveness.
D.2 Ethical considerations